A​ monopolist's maximized rate of economic profits is ​$1 comma 800 per week. Its weekly output is 900 ​units, and at this output​ rate, the​ firm's marginal cost is ​$34 per unit. The price at which it sells each unit is ​$46 per unit.

At these profit and output​ rates, the​ firm's average total cost is ​$ nothing. ​(Enter your response as a whole​ number.) At these profit and output​ rates, the​ firm's marginal revenue is ​$ nothing. ​(Enter your response as a whole​ number.)
Answer #1

a monopolist's maximized rate of economic profits = 1800

MC = 34

Q = 900

P = 46

Profit = (P-ATC)*Q

1800 = (46-ATC)*900

(46-ATC) = 1800 / 900 = 2

ATC = 46 - 2 = 44

Since firm is maximizing profit, means MR = MC

So MR = MC = 34